# Investigating Data Pruning for Pretraining Biological Foundation Models at Scale

### v1

@@ 2.2 Biological Sequence Representation
Learning effective representations of biological sequences—RNA, DNA, and proteins—is fundamental to a wide range of downstream tasks, including structure prediction, function annotation, and biomolecular interaction modeling (Shen et al. 2024). Due to the intrinsic complexity of biological macromolecules, such as RNA’s hierarchical structure and protein folding patterns, traditional approaches often relied on hand-crafted features or shallow learning techniques specific to individual tasks. Recent advances in biological foundation models (BioFMs) (Chen et al. 2022; Hayes et al. 2025; Brixi et al. 2025) have demonstrated the effectiveness of large-scale self-supervised learning, particularly masked language modeling, in capturing intricate biological semantics directly from raw sequences. For RNA, models, such as RNA-FM (Chen et al. 2022), have achieved strong transfer performance across diverse tasks, including RNA type classification (Amin, McGrath, and Chen 2019), CRISPR-Cas efficiency prediction (Chuai et al. 2018), and RNA-binding protein (RBP) interaction prediction (Xu et al. 2023; Zhu et al. 2023). On the protein side, ESM (Hayes et al. 2025) has shown promising results on protein function prediction (Sarkisyan et al. 2016), protein structure prediction (Klausen et al. 2019), and interaction prediction (Guo et al. 2008). Detailed descriptions of RNA-FM and ESM models, including their architectures, are provided in Appendix A.2. Despite their success, BioFMs suffer from high training costs and limited reproducibility, which further motivates the need for post-hoc data-efficient approaches, such as data pruning, that can identify informative training subsets without full retraining access.
@@ 3.2 Subset-Based Scalable Influence Estimation
Previous studies (Kim, Kim, and Yang 2023) have shown that second-order approximations yield more accurate estimates of loss changes. Therefore, we adopt a second-order Taylor expansion to estimate the loss change:
@@
To further reduce computational cost, we follow the practice commonly adopted in conjugate gradient methods (Roux, Manzagol, and Bengio 2007; Schaul, Zhang, and LeCun 2013; Martens and Grosse 2015) and adaptive optimizers such as Adam (Kingma and Ba 2015), where the curvature matrix is approximated by its diagonal. In this spirit, we apply a diagonal approximation to the empirical Fisher matrix, which yields:
@@ 4.2 Downstream Tasks and Evaluation Metrics
  • Function Prediction: (1) RNA Type Classification (TypeCls) (Amin, McGrath, and Chen 2019), evaluated using accuracy and F1 score; (2) RNA Modification Prediction (Modif) (Duan, Wang, and Jia 2019), evaluated using AUC.
@@ Acknowledgments
This work was supported by the Major Project of Guangzhou National Laboratory (Grant No. GZNL2024A01003, GZNL2023A02007, GZNL2025C02028 to Jiao Yuan), National Natural Science Foundation of China (Grant No. 32400547 to Jiao Yuan), Pearl River Talent Recruitment Program (2023QN10Y296 to Jiao Yuan), Guangzhou Young Top Talent Program, National Key R&D Program of China (2023YFF1204701 to Jiao Yuan), the Chinese University of Hong Kong (CUHK; award numbers 4937025, 4937026, 5501517 and 5501329 to Yu Li), Shenzhen Medical Research Fund (Grant No. A2503002 to Yu Li), the IdeaBooster Fund (IDBF23ENG05 and IDBF24ENG06 to Yu Li), partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region (Hong Kong SAR), China (project no. CUHK 24204023 to Yu Li), a grant from the Innovation and Technology Commission of the Hong Kong SAR, China (project no. GHP/065/21SZ and ITS/247/23FP to Yu Li), and the Research Matching Grant Scheme at CUHK (award numbers 8601603 and 8601663 to Yu Li) from the Research Grants Council, Hong Kong SAR, China.
@@ A.2 Model Details of Biological Foundation Models
(RNA Foundation Model) (Chen et al. 2022) is a self-supervised deep learning framework based on the BERT architecture(Devlin et al. 2018), employing 12 bidirectional Transformer encoder layers to extract biologically meaningful representations from RNA sequences. It is pretrained on 23 million non-coding RNA sequences using a masked language modeling (MLM) objective. Following pretraining, RNA-FM can be fine-tuned for various downstream tasks, including secondary structure prediction (Danaee et al. 2018), contact map prediction (Leontis and Zirbel 2012), RBP interaction modeling (Xu et al. 2023; Zhu et al. 2023), RNA type classification (Amin, McGrath, and Chen 2019), and CRISPR-Cas efficiency prediction (Chuai et al. 2018).
@@ B.1 RNA Downstream Tasks
This task is to predict each RNA sequence into one of 10 functional RNA types (e.g., rRNA, tRNA, snoRNA) (Amin, McGrath, and Chen 2019). We evaluate model performance using accuracy (overall correctness) and F1 score (robustness under class imbalance). A linear classification head is appended to the mean-pooled representation from the RNA-FM encoder, and all model parameters are fine-tuned using cross-entropy loss. The model is trained for 30 epochs with a batch size of 32 using the AdamW optimizer. We set the learning rate to $1\times 10^{-4}$ and apply a weight decay of 0.01 throughout training.
@@
This task is to predict the presence of 12 common RNA modifications for a given input sequence (Duan, Wang, and Jia 2019). Model performance is evaluated using the mean AUC (area under the ROC curve). We append a sigmoid-activated linear classification head on top of the mean pooled representation from the RNA-FM encoder and fine-tune all model parameters using binary cross-entropy loss. The model is trained for 30 epochs with a batch size of 32 using the AdamW optimizer. We set the learning rate to $5\times 10^{-5}$, and apply a weight decay of 0.01 throughout training.

### v2

@@ 2.2 Biological Sequence Representation
Learning effective representations of biological sequences—RNA, DNA, and proteins—is fundamental to a wide range of downstream tasks, including structure prediction, function annotation, and biomolecular interaction modeling (Shen et al. 2024). Due to the intrinsic complexity of biological macromolecules, such as RNA’s hierarchical structure and protein folding patterns, traditional approaches often relied on hand-crafted features or shallow learning techniques specific to individual tasks. Recent advances in biological foundation models (BioFMs) (Chen et al. 2022; Hayes et al. 2025; Brixi et al. 2025) have demonstrated the effectiveness of large-scale self-supervised learning, particularly masked language modeling, in capturing intricate biological semantics directly from raw sequences. For RNA, models, such as RNA-FM (Chen et al. 2022), have achieved strong transfer performance across diverse tasks, including RNA type classification (Amin et al. 2019), CRISPR-Cas efficiency prediction (Chuai et al. 2018), and RNA-binding protein (RBP) interaction prediction (Xu et al. 2023; Zhu et al. 2023). On the protein side, ESM (Hayes et al. 2025) has shown promising results on protein function prediction (Sarkisyan et al. 2016), protein structure prediction (Klausen et al. 2019), and interaction prediction (Guo et al. 2008). Detailed descriptions of RNA-FM and ESM models, including their architectures, are provided in Appendix A.2. Despite their success, BioFMs suffer from high training costs and limited reproducibility, which further motivates the need for post-hoc data-efficient approaches, such as data pruning, that can identify informative training subsets without full retraining access.
@@ 3.2 Subset-Based Scalable Influence Estimation
Previous studies (Kim et al. 2023) have shown that second-order approximations yield more accurate estimates of loss changes. Therefore, we adopt a second-order Taylor expansion to estimate the loss change:
@@
To further reduce computational cost, we follow the practice commonly adopted in conjugate gradient methods (Roux et al. 2007; Schaul et al. 2013; Martens and Grosse 2015) and adaptive optimizers such as Adam (Kingma and Ba 2015), where the curvature matrix is approximated by its diagonal. In this spirit, we apply a diagonal approximation to the empirical Fisher matrix, which yields:
@@ 4.2 Downstream Tasks and Evaluation Metrics
  • Function Prediction: (1) RNA Type Classification (TypeCls) (Amin et al. 2019), evaluated using accuracy and F1 score; (2) RNA Modification Prediction (Modif) (Duan et al. 2019), evaluated using AUC.
@@ Acknowledgments
This work was supported by Shenzhen Medical Research Fund (Grant No. A2503002 to Yu Li), the Major Project of Guangzhou National Laboratory (Grant No. GZNL2024A01003, GZNL2023A02007, GZNL2025C02028 to Jiao Yuan), the National Natural Science Foundation of China (Grant No. 32400547 to Jiao Yuan and Grant No. 62472097 to Weizhong Zhang), High-Quality Development Project of Shanghai Municipal Commission of Economy and Informatization (Grant No. 2024-GZL-RGZN-02010 to Weizhong Zhang), AI for Science Foundation of Fudan University (FudanX24AI028 to Weizhong Zhang), Pearl River Talent Recruitment Program (2023QN10Y296 to Jiao Yuan), Guangzhou Young Top Talent Program, National Key R&D Program of China (2023YFF1204701 to Jiao Yuan), the Chinese University of Hong Kong (CUHK; award numbers 4937025, 4937026, 5501517 and 5501329 to Yu Li), the IdeaBooster Fund (IDBF23ENG05 and IDBF24ENG06 to Yu Li), partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region (Hong Kong SAR), China (project no. CUHK 24204023 to Yu Li), a grant from the Innovation and Technology Commission of the Hong Kong SAR, China (project no. GHP/065/21SZ and ITS/247/23FP to Yu Li), and the Research Matching Grant Scheme at CUHK (award numbers 8601603 and 8601663 to Yu Li) from the Research Grants Council, Hong Kong SAR, China.
@@ A.2 Model Details of Biological Foundation Models
(RNA Foundation Model) (Chen et al. 2022) is a self-supervised deep learning framework based on the BERT architecture(Devlin et al. 2018), employing 12 bidirectional Transformer encoder layers to extract biologically meaningful representations from RNA sequences. It is pretrained on 23 million non-coding RNA sequences using a masked language modeling (MLM) objective. Following pretraining, RNA-FM can be fine-tuned for various downstream tasks, including secondary structure prediction (Danaee et al. 2018), contact map prediction (Leontis and Zirbel 2012), RBP interaction modeling (Xu et al. 2023; Zhu et al. 2023), RNA type classification (Amin et al. 2019), and CRISPR-Cas efficiency prediction (Chuai et al. 2018).
@@ B.1 RNA Downstream Tasks
This task is to predict each RNA sequence into one of 10 functional RNA types (e.g., rRNA, tRNA, snoRNA) (Amin et al. 2019). We evaluate model performance using accuracy (overall correctness) and F1 score (robustness under class imbalance). A linear classification head is appended to the mean-pooled representation from the RNA-FM encoder, and all model parameters are fine-tuned using cross-entropy loss. The model is trained for 30 epochs with a batch size of 32 using the AdamW optimizer. We set the learning rate to $1\times 10^{-4}$ and apply a weight decay of 0.01 throughout training.
@@
This task is to predict the presence of 12 common RNA modifications for a given input sequence (Duan et al. 2019). Model performance is evaluated using the mean AUC (area under the ROC curve). We append a sigmoid-activated linear classification head on top of the mean pooled representation from the RNA-FM encoder and fine-tune all model parameters using binary cross-entropy loss. The model is trained for 30 epochs with a batch size of 32 using the AdamW optimizer. We set the learning rate to $5\times 10^{-5}$, and apply a weight decay of 0.01 throughout training.